OpenStudy (leyla.tas123):

what is the length of each side of a square if the A=9 m^2

jhonyy9 (jhonyy9):

do you know formula for area of a square ?

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

how you get area of a square ?

jhonyy9 (jhonyy9):

@OtherWorldly

jhonyy9 (jhonyy9):

area of a square equal side time side so A = s*s A = s^2

jhonyy9 (jhonyy9):

can you calculi from this the length of a side ?

jhonyy9 (jhonyy9):

9 = s^2 s = sqrt(9) s = ?

OpenStudy (mathmale):

I am assuming that this equation is correct: A=9 m^2 Our task is to solve this for the length of one side of the square. Start by dividing both sides of this equation by 9, to isolate "m^2" on the right. Then find the positive square root of both sides. Your answer?